OpenStudy (anonymous):

What are the complex solutions of x^2+10=2x? A x=1+3i, x=1-3i B x=3+i, x=3-i C x=-1+3i, x=-1-3i D x=-5, x=2

OpenStudy (anonymous):

1+3i and 1-3i are the complex roots of this equation

OpenStudy (anonymous):

Could you show me how you got that please?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2-2x+10=0 now compare with ax^2+bx+c here a=1 b=-2 and c=10 now find the discriminant d which is b^2-4ac u will get it as -36 now from this we can say the equation has two complex roots which are \[-b {\pm}\sqrt{d}/2a\] so just substitute the value of d ,a ,b in the equation and u will get ur answer

OpenStudy (anonymous):

Okay, on my other question you answered about the distance, could you set up that equation for me. I was having a little trouble understanding it

OpenStudy (anonymous):

yes during the first trip let the planes velocity due to its engine be v but as the wind is in opposite direction the actual velocity due which the plane is moving is v-50 so we can say v-50=x/6 where x is the distance between the two cities now in the return trip the velocity of plane is v but the wind velocity is in assisting condition so the actual velocity becomes v+50 so v+50=x/4, now from these eq.. calculate v and then x